Case The city of Flinth, Arizona needs a new water delivery system from Lake Not There Yet. The two options proposed by the cQuestion Three Your recommendation to the city council is: [Select]
Answer #1

Equivalent annual cost of option one = 3500000 * (A/P, 5.5%,30) + 15000

= 3500000 * 0.068805 + 15000

= 255818.86 ~ 255819 (Nearest Dollar)

For option two

Present value of geometric series = A * [1 - (1+g)^n / (1+i)^n] / (i-g)

We need to find present value of power cost for 20 yrs

Present value of power cost =  30000 * [1 - (1+0.065)^20 / (1+0.055)^20] / (0.055-0.065)

=  30000 * [1 - (1.065)^20 / (1.055)^20] / (-0.01)

= 30000 * 20.76552198

= 622965.66

Equivalent annual cost of option two = (2500000 + 200000 * (P/F, 5.5%,8) + 622965.66) *(A/P, 5.5%,20) + 25000

= (2500000 + 200000 * 0.651599 + 622965.66) * 0.083679 + 25000

= 297232.75 ~ 297233 (Nearest Dollar)

As Equivalent annual cost of option one is lower it should be selected